# Spectral Convolutional Conditional Neural Processes

### v6

@@ 2.2 Neural Processes
Let $\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$ denote the space of probability measures over measurable functions $f:\mathcal{X}\to\mathcal{Y}$, which we interpret as stochastic processes indexed by $\mathcal{X}$. We assume an unknown data-generating process $\mu\in\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, from which latent functions are drawn. A task consists of a finite collection of noisy input–output observations generated from a realization $f\sim\mu$, partitioned into a context set and a query set:
@@
where $\epsilon_{c,k},\epsilon_{q,l}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma_{0}^{2})$ and $\sigma_{0}>0$ denotes the observation noise scale.
@@
$\varphi_{e}[\mathcal{D}_{c}](x)=\sum_{(x_{c},y_{c})\in\mathcal{D}_{c}}\!\!\!\!\phi(y_{c})\,\psi_{e}(x-x_{c}).$ (3)
@@ 4 Experiments
We evaluate our framework on four regression benchmarks and compare its performance against several members of CNPs family. Specifically, we include the original CNP (Garnelo et al., 2018a), the Attentive CNP (AttCNP; Kim et al. (2019)), the Convolutional CNP (ConvCNP; Gordon et al. (2019)), the diagonal variant of the Transformer Neural Process (TNP; Nguyen and Grover (2022)), and the Translation-Equivariant Transformer Neural Process (TETNP; Ashman et al. (2024a)). For each experimental setting, all models are trained using four random seeds. Performance is evaluated using the final aggregated log-likelihood and root-mean-squared error (RMSE), and we report the mean $\pm$ standard deviation across runs. Additional details regarding datasets, architectures, and training procedures are provided in Appendix B. Our implementation and experimental code are publicly available at https://github.com/peiman-m/SConvCNP.
@@ 4.1 Synthetic 1-D Regression
CNP AttCNP TNP TETNP ConvCNP SConvCNP
@@
We begin by evaluating models on four synthetic benchmarks generated from distinct stochastic processes: a GP with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. For each benchmark, the parameters of the generative process—kernel hyperparameters for the GPs, frequency and direction for the sawtooth generator, and frequency and duty cycle for the square-wave generator—are sampled randomly. Complete experimental details are provided in Appendix B.1. Table 1 reports the average evaluation metrics over 1,000 test batches, each containing 16 tasks (each task defined in equation 2). For every task—independent of batch—both $x_{c}$ and $x_{q}$ are sampled independently from $\mathcal{U}[-3,3)$. For each test batch, the number of context points, shared by all tasks within that batch, is drawn independently according to $n_{c}\sim\mathcal{U}[5,25)$. The number of query points, however, is fixed at $n_{q}=256$ for all the test tasks. As shown, SConvCNP consistently outperforms or closely matches strong baselines, including TETNP and ConvCNP, which represent the current state of the art. Figure 1 provides qualitative comparisons of predictive maps produced by SConvCNP, ConvCNP, and TETNP.
For the sawtooth-wave benchmark, we were unable to successfully train either TNP or TETNP. Across the configurations we attempted—including increased model capacity and multiple randomized initializations—their predictions consistently collapsed to zero. We hypothesize that this failure mode stems from spectral bias, the tendency of neural networks to favor low-frequency structure over high-frequency components (Rahaman et al., 2019; Ronen et al., 2019; Basri et al., 2020; Tancik et al., 2020; Fridovich-Keil et al., 2022). Recent findings by Vasudeva et al. (2025) further suggest that this bias is exacerbated in transformer architectures compared with convolutional networks. This interpretation aligns with observations by Nguyen and Grover (2022), who originally introduced TNPs. They report degraded performance on GP samples drawn from a periodic kernel—reflected in poor log-likelihood scores—despite strong results on Matérn-kernel tasks. Interestingly, our experiments on periodic-kernel GP tasks (Table 1) do not replicate this limitation. A key distinction is that we employ the Efficient Query TNP (EQTNP) architecture of Feng et al. (2022), rather than the original TNP design (Nguyen and Grover, 2022) (see Appendix B.1.1 for architectural details). Finally, although sawtooth and square-wave functions share similar discontinuity and high-frequency characteristics, Table 1 shows that TNP and TETNP do not collapse when trained on square-wave signals. Investigating the source of this discrepancy falls beyond the scope of this work but represents an intriguing direction for future study.
@@ 4.2 Predator–Prey Dynamics
Table 2 reports average evaluation metrics over 1,000 test batches, each containing 16 tasks. For each batch, the number of context points—shared across all tasks—is sampled as $n_{c}\sim\mathcal{U}[5,25)$, and the number of query points is fixed at $n_{q}=256$. SConvCNP achieves log-likelihood performance comparable to ConvCNP, and both are competitive with TETNP. In terms of RMSE, SConvCNP matches TETNP, indicating that its predictive uncertainty intervals are slightly wider. Qualitative comparisons of predictive maps for the top three models appear in Figure 2.
@@
CNP AttCNP TNP TETNP ConvCNP SConvCNP
@@ 4.3 Traffic Flow
CNP AttCNP TNP TETNP ConvCNP SConvCNP
@@ 4.4 Image Completion
For our final experiment, we evaluate model performance on an image-completion task formulated as a spatial regression problem, where the model maps 2D pixel coordinates to their corresponding intensity values. We use images from the Describable Textures Dataset (DTD; Cimpoi et al. (2014)), and construct each task from a processed $64\times 64$ subsampled crop of an original image. Because each task contains a large number of context and query pixels, we were unable to fit TETNP within our computational budget—even after reducing its size—so we exclude it from this experiment. Additional experimental details are provided in Appendix B.4. Table 4 reports results averaged over 1,880 test tasks, evaluated in batches of 16. For each batch, the number of context points—shared across all tasks—is sampled as $n_{c}\sim\mathcal{U}[5,1024)$; all remaining pixels serve as query points, so that $n_{q}=4096-n_{c}$. Among the baselines, SConvCNP achieves the highest log-likelihood and the second-lowest RMSE. Figure 4 presents qualitative comparisons of the predictive means produced by the top three models.
@@ B.1.1 Model Architectures
Translation-Equivariant Transformer Neural Process (TETNP).
The token construction in the TETNP closely follows that of the TNP, with a crucial modification to ensure translation equivariance: input locations are excluded from the token representations. Specifically, for each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$, we construct the token
@@
The TETNP replaces the standard multi-head attention mechanism with the translation-equivariant attention proposed by Ashman et al. [2024a]. For each attention head, we first compute the matrix of pairwise scaled dot products between token embeddings, analogous to conventional attention but independent of absolute input locations. In parallel, we compute the matrix of pairwise differences between token locations. We then concatenate the scaled dot-product scores from all heads with the corresponding pairwise location differences. This augmented pairwise similarity representation is processed by an MLP—acting as an implicit kernel—with two hidden layers of width 256 and an output dimension equal to the number of attention heads, yielding the final translation-equivariant attention logits.
@@
CNP AttCNP TNP TETNP ConvCNP SConvCNP
@@
CNP AttCNP TNP TETNP ConvCNP SConvCNP
@@ B.1.2 Data and Experimental Setup
Process Parameters.
For each task family, hyperparameters are sampled independently at the task level:
  • GP with Matérn–5/2 kernel. The scaled distance is
    $d=\sqrt{(x-x^{\prime})^{\top}\Lambda^{-2}(x-x^{\prime})},\qquad\Lambda=\lambda I,\quad\lambda\sim\mathcal{U}[0.25,1).$
    The kernel is
    $k_{m2.5}(x,x^{\prime})=\frac{2^{-1.5}}{\Gamma(2.5)}\,(\sqrt{5}\,d)^{2.5}K_{2.5}(\sqrt{5}\,d),$
    where $K_{2.5}$ is a modified Bessel function of the second kind. Observation noise has variance $\sigma_{0}^{2}=0.01$.
  • GP with periodic kernel. The kernel is parameterized by period $\rho\sim\mathcal{U}[0.5,2)$ and lengthscale $\lambda\sim\mathcal{U}[0.25,1)$:
    $k_{p}(x,x^{\prime})=\exp\!\left(-2\sum_{i=1}^{d}\frac{\sin^{2}\!\big(\frac{\pi}{\rho}(x_{i}-x_{i}^{\prime})\big)}{\lambda}\right),$
    with observation noise variance $\sigma_{0}^{2}=0.01$.
  • Sawtooth wave. Piecewise-linear periodic functions with frequency $\xi\sim\mathcal{U}(0.5,5)$, direction $u\in\{+1,-1\}$, and offset $c\sim\mathcal{U}[0,1)$. Noise follows $\mathcal{N}(0,0.0025)$.
  • Square wave. Discontinuous periodic functions with frequency $\xi\sim\mathcal{U}(0.5,5)$, duty cycle $D\sim\mathcal{U}[0.25,0.75)$, and offset $c\sim\mathcal{U}[0,1)$. Noise follows $\mathcal{N}(0,0.0025)$.
Table 7: Synthetic generative processes used in 1D regression experiments.
Process $f(x)$
GP (Matérn 5/2) $f\sim\mathcal{GP}(0,k_{m2.5}+\sigma_{0}^{2}I)$
GP (Periodic) $f\sim\mathcal{GP}(0,k_{p}+\sigma_{0}^{2}I)$
Sawtooth $f(x)=2\left(\left(\xi\left(u^{\top}x-c\right)\right)\bmod 1\right)+\sigma_{0}(x),\,\,\sigma_{0}(x)\sim\mathcal{N}(0,\sigma_{0}^{2})$
Square Wave $f(x)=2\,\mathds{1}_{\left\{\left(\left(\xi x-c\right)\bmod 1\right)<D\right\}}-1+\sigma_{0}(x),\,\,\sigma_{0}(x)\sim\mathcal{N}(0,\sigma_{0}^{2})$

### v7

@@ 2.2 Neural Processes
Let $\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$ denote the space of probability measures over measurable functions $f:\mathcal{X}\to\mathcal{Y}$, which we interpret as stochastic processes indexed by $\mathcal{X}$. We assume an unknown data-generating process $\mu\in\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, from which latent functions are drawn. A task consists of a finite number of noisy input–output observations generated from a realization $f\sim\mu$, partitioned into a context set and a query set:
@@
where $\epsilon_{c,k},\epsilon_{q,l}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma_{0}^{2})$ and $\sigma_{0}>0$ denotes the observation noise standard deviation.
@@
$\varphi_{e}[\mathcal{D}_{c}](x)=\mkern-15.0mu\sum_{(x_{c},y_{c})\in\mathcal{D}_{c}}\mkern-15.0mu\phi(y_{c})\,\psi_{e}(x-x_{c}).$ (3)
@@ 4 Experiments
We evaluate our framework on four regression benchmarks and compare its performance against several members of CNPs family. Specifically, we include the original CNP (Garnelo et al., 2018a), the Attentive CNP (AttCNP; Kim et al. (2019)), the Convolutional CNP (ConvCNP; Gordon et al. (2019)), the diagonal variant of the Transformer Neural Process (TNP; Nguyen and Grover (2022)), and the Translation-Equivariant Transformer Neural Process (TE-TNP; Ashman et al. (2024a)). For each experimental setting, all models are trained using four random seeds. Performance is evaluated using the final aggregated log-likelihood and root-mean-squared error (RMSE), and we report the mean $\pm$ standard deviation across runs. Additional details regarding datasets, architectures, and training procedures are provided in Appendix B. Our implementation and experimental code are publicly available at https://github.com/peiman-m/SConvCNP.
@@ 4.1 Synthetic 1-D Regression
CNP AttCNP TNP TE-TNP ConvCNP SConvCNP
@@
We begin by evaluating models on four synthetic benchmarks generated from distinct stochastic processes: a GP with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. For each benchmark, the parameters of the generative process—kernel hyperparameters for the GPs, frequency and direction for the sawtooth generator, and frequency and duty cycle for the square-wave generator—are sampled randomly. Complete experimental details are provided in Appendix B.1. Table 1 reports the average evaluation metrics over 1,000 test batches, each containing 16 tasks (each task defined in equation 2). For every task—independent of batch—both $x_{c}$ and $x_{q}$ are sampled independently from $\mathcal{U}[-3,3)$. For each test batch, the number of context points, shared by all tasks within that batch, is drawn independently according to $n_{c}\sim\mathcal{U}[5,25)$. The number of query points, however, is fixed at $n_{q}=256$ for all the test tasks. As shown, SConvCNP consistently outperforms or closely matches strong baselines, including TE-TNP and ConvCNP, which represent the current state of the art. Figure 1 provides qualitative comparisons of predictive maps produced by SConvCNP, ConvCNP, and TE-TNP.
For the sawtooth-wave benchmark, we were unable to successfully train either TNP or TE-TNP. Across the configurations we attempted—including increased model capacity and multiple randomized initializations—their predictions consistently collapsed to zero. We hypothesize that this failure mode stems from spectral bias, the tendency of neural networks to favor low-frequency structure over high-frequency components (Rahaman et al., 2019; Ronen et al., 2019; Basri et al., 2020; Tancik et al., 2020; Fridovich-Keil et al., 2022). Recent findings by Vasudeva et al. (2025) further suggest that this bias is exacerbated in transformer architectures compared with convolutional networks. This interpretation aligns with observations by Nguyen and Grover (2022), who originally introduced TNPs. They report degraded performance on GP samples drawn from a periodic kernel—reflected in poor log-likelihood scores—despite strong results on Matérn-kernel tasks. Interestingly, our experiments on periodic-kernel GP tasks (Table 1) do not replicate this limitation. A key distinction is that we employ the Efficient Query TNP (EQTNP) architecture of Feng et al. (2022), rather than the original TNP design (Nguyen and Grover, 2022) (see Appendix B.1.1 for architectural details). Finally, although sawtooth and square-wave functions share similar discontinuity and high-frequency characteristics, Table 1 shows that TNP and TE-TNP do not collapse when trained on square-wave signals. Investigating the source of this discrepancy falls beyond the scope of this work but represents an intriguing direction for future study.
@@ 4.2 Predator–Prey Dynamics
Table 2 reports average evaluation metrics over 1,000 test batches, each containing 16 tasks. For each batch, the number of context points—shared across all tasks—is sampled as $n_{c}\sim\mathcal{U}[5,25)$, and the number of query points is fixed at $n_{q}=256$. SConvCNP achieves log-likelihood performance comparable to ConvCNP, and both are competitive with TE-TNP. In terms of RMSE, SConvCNP matches TE-TNP, indicating that its predictive uncertainty intervals are slightly wider. Qualitative comparisons of predictive maps for the top three models appear in Figure 2.
@@
CNP AttCNP TNP TE-TNP ConvCNP SConvCNP
@@ 4.3 Traffic Flow
CNP AttCNP TNP TE-TNP ConvCNP SConvCNP
@@ 4.4 Image Completion
For our final experiment, we evaluate model performance on an image-completion task formulated as a spatial regression problem, where the model maps 2D pixel coordinates to their corresponding intensity values. We use images from the Describable Textures Dataset (DTD; Cimpoi et al. (2014)), and construct each task from a processed $64\times 64$ subsampled crop of an original image. Because each task contains a large number of context and query pixels, we were unable to fit TE-TNP within our computational budget—even after reducing its size—so we exclude it from this experiment. Additional experimental details are provided in Appendix B.4. Table 4 reports results averaged over 1,880 test tasks, evaluated in batches of 16. For each batch, the number of context points—shared across all tasks—is sampled as $n_{c}\sim\mathcal{U}[5,1024)$; all remaining pixels serve as query points, so that $n_{q}=4096-n_{c}$. Among the baselines, SConvCNP achieves the highest log-likelihood and the second-lowest RMSE. Figure 4 presents qualitative comparisons of the predictive means produced by the top three models.
@@ B.1.1 Model Architectures
Translation-Equivariant Transformer Neural Process (TE-TNP).
The token construction in the TE-TNP closely follows that of the TNP, with a crucial modification to ensure translation equivariance: input locations are excluded from the token representations. Specifically, for each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$, we construct the token
@@
The TE-TNP replaces the standard multi-head attention mechanism with the translation-equivariant attention proposed by Ashman et al. [2024a]. For each attention head, we first compute the matrix of pairwise scaled dot products between token embeddings, analogous to conventional attention but independent of absolute input locations. In parallel, we compute the matrix of pairwise differences between token locations. We then concatenate the scaled dot-product scores from all heads with the corresponding pairwise location differences. This augmented pairwise similarity representation is processed by an MLP—acting as an implicit kernel—with two hidden layers of width 256 and an output dimension equal to the number of attention heads, yielding the final translation-equivariant attention logits.
@@
CNP AttCNP TNP TE-TNP ConvCNP SConvCNP
@@
CNP AttCNP TNP TE-TNP ConvCNP SConvCNP
@@ B.1.2 Data and Experimental Setup
Generative Processes.
For each task family, process-specific hyperparameters are sampled independently at the task level.
  • GP with Matérn–5/2 kernel. Functions are sampled from $f\sim\mathcal{GP}(0,k_{\text{m5/2}}+\sigma_{0}^{2}I)$, where
    $k_{\text{m5/2}}(x,x^{\prime})=\frac{2^{-1.5}}{\Gamma(2.5)}\,(\frac{\sqrt{5}}{\lambda}|x-x^{\prime}|)^{2.5}K_{2.5}(\frac{\sqrt{5}}{\lambda}|x-x^{\prime}|)$
    and $K_{2.5}$ is the modified Bessel function of the second kind. The lengthscale is sampled as $\lambda\sim\mathcal{U}[0.25,1)$. Observation noise standard deviation is $\sigma_{0}=0.1$.
  • GP with periodic kernel. Functions are sampled from $f\sim\mathcal{GP}(0,k_{\text{p}}+\sigma_{0}^{2}I)$, where
    $k_{\text{p}}(x,x^{\prime})=\exp\!\left(-\frac{2\sin^{2}(\pi|x-x^{\prime}|/\rho)}{\lambda^{2}}\right)$
    with period $\rho\sim\mathcal{U}[0.5,2)$ and lengthscale $\lambda\sim\mathcal{U}[0.25,1)$. Observation noise standard deviation is $\sigma_{0}=0.1$.
  • Sawtooth wave. Functions are sampled from $f\sim\mathcal{GP}(m_{\text{saw}},\sigma_{0}^{2}I)$, where the mean function is
    $m_{\text{saw}}(x)=2\left(\left(\xi\left(ux-c\right)\right)\bmod 1\right)-1$
    with frequency $\xi\sim\mathcal{U}[0.5,5)$, direction $u\in\{+1,-1\}$ (sampled uniformly), and phase offset $c\sim\mathcal{U}[0,1)$. Observation noise standard deviation is $\sigma_{0}=0.05$.
  • Square wave. Functions are sampled from $f\sim\mathcal{GP}(m_{\text{sq}},\sigma_{0}^{2}I)$, where the mean function is
    $m_{\text{sq}}(x)=2\,\mathds{1}_{\{((\xi x-c)\bmod 1)<D\}}-1$
    with frequency $\xi\sim\mathcal{U}[0.5,5)$, duty cycle $D\sim\mathcal{U}[0.25,0.75)$, and phase offset $c\sim\mathcal{U}[0,1)$. Observation noise standard deviation is $\sigma_{0}=0.05$.